5. In testing a new drug, researchers found that 5% of all patients using it will have a mild side effect. A random sample of 12 patients using the drug is selected. Find the probability that:

(A) exactly three will have this mild side effect
(B) at least two will have this mild side effect.

Answer by stanbon(57337)

In testing a new drug, researchers found that 5% of all patients using it will have a mild side effect. A random sample of 12 patients using the drug is selected.
-------------------
Binomial Problem with n = 12 ; p = 0.05
-----------------------------
Find the probability that:
(A) exactly three will have this mild side effect
P(x=3) = 0.05^3 = 0.125
--------------------------------------------------
(B) at least two will have this mild side effect.
P(2<=x<=12) = 1 - P(0<= x <=1) = 0.1184
====================================================
Cheers,
Stan H.